# Cooperative Multi-Agent Planning with Adaptive Skill Synthesis

### v1

@@ Abstract
Despite much progress in training distributed artificial intelligence (AI), building cooperative multi-agent systems with multi-agent reinforcement learning (MARL) faces challenges in sample efficiency, interpretability, and transferability. Unlike traditional learning-based methods that require extensive interaction with the environment, large language models (LLMs) demonstrate remarkable capabilities in zero-shot planning and complex reasoning. However, existing LLM-based approaches heavily rely on text-based observations and struggle with the non-Markovian nature of multi-agent interactions under partial observability. We present COMPASS, a novel multi-agent architecture that integrates vision-language models (VLMs) with a dynamic skill library and structured communication for decentralized closed-loop decision-making. The skill library, bootstrapped from demonstrations, evolves via planner-guided tasks to enable adaptive strategies. COMPASS propagates entity information through multi-hop communication under partial observability. Evaluations on the improved StarCraft Multi-Agent Challenge (SMACv2) demonstrate COMPASS achieves up to 30% higher win rates than state-of-the-art MARL algorithms in symmetric scenarios.
Machine Learning, ICML
@@ 1 Introduction
A major long-term goal for the field of cooperative multi-agent systems (MAS), e.g. multi-robot control (Gao et al., 2024; Feng et al., 2024), power management (Monroc et al., 2024), and multi-agent games (Samvelyan et al., 2019; Kurach et al., 2020), is to build a protocol of collaboration among the agents. Multi-agent reinforcement learning (MARL), proven as an advanced paradigm of distributed artificial intelligence (AI), holds promise for discovering collective behavior from interactions. One line of works follows centralized training decentralized execution (CTDE) paradigm (Rashid et al., 2020; Yu et al., 2022; Liu et al., 2024; Lowe et al., 2017; Zhu et al., 2025; Li et al., 2024b). CTDE assumes a central controller that exploits global information, while the individual policies are designed to allow for decentralized execution. However, in many real-world scenarios, the central controller becomes unfeasible due to the communication overhead that exponentially scales with the number of agents, thereby compromising the scalability of MAS. In contrast, the decentralized training decentralized execution paradigm (DTDE) discards this assumption and is scalable to large-scale systems (Su & Lu, 2024; Zhang et al., 2018; Ma et al., 2024a). However, DTDE requires complicated learning and planning under uncertainty, as partial observability magnifies the discrepancy between each agent’s local observation and global information. Although much progress has been made, MARL suffers from compromised sample efficiency, interpretability, and transferability.
@@
The emergence of Large Language Models (LLMs) has revitalized this field. LLM-based multi-agents have been proposed to leverage their remarkable capacity to perform task-oriented collective behaviors (Mandi et al., 2024; Zhang et al., 2024; Gong et al., 2024; Zhang et al., 2025; Nayak et al., 2025). The LLMs are used for high-level planning to generate centralized (Gong et al., 2024; Nayak et al., 2025; Deng et al., 2024) or decentralized plans (Mandi et al., 2024; Zhang et al., 2024), often adopting a hierarchical decision-making structure in conjunction with a pre-defined low-level controller. While these methods have succeeded in a set of multi-agent problems including Overcooked-AI (Carroll et al., 2019), SMAC (Samvelyan et al., 2019), and VirtualHome (Puig et al., 2018), heavy reliance on text-based observation prevents them from learning from multi-modal information. Moreover, they ignore the non-Markovian nature of MAS, where learning and planning necessitate a decentralized closed-loop solution.
@@
In this paper, we mitigate the existing research limitations and advance general decision-making for cooperative multi-agent systems. At a high level, COMPASS combines a vision language models (VLMs)-based planner with a dynamic skill library for storing and retrieving complex behaviors, along with a structured communication protocol. A diagram of COMPASS is provided in Figure. 1. Inspired by Cradle (Tan et al., 2024), the VLM-based planner perceives the visual and textual observation and suggests the most suitable executable code from the skill library. We adopt the code-as-policy paradigm (Wang et al., 2024d) instead of task-specific primitive actions, as it constrains generalizability and fails to fully leverage foundation models’ extensive world knowledge and sophisticated reasoning capabilities.
@@
Traditional open-loop methods struggle to produce effective plans that adapt to dynamics in stochastic, partially observable environments. To address this challenge, the VLM-based planer attempts to solve challenging and ambiguous final tasks, such as ”Defeat all enemy units in the StarCraft multi-agent combat scenario while coordinating with allied units”, by progressively proposing a sequence of clear, manageable sub-tasks while incorporating environmental feedback and task progress. COMPASS generates Python scripts through LLMs as semantic-level skills to accomplish sub-tasks, incrementally building a skill library throughout the task progress. Each skill is indexed through its documentation embeddings, enabling retrieval based on task-skill relevance. However, developing the skill library from scratch requires extensive exploration to discover viable strategies. In contrast, we pre-collect demonstration videos and introduce a ”warm start” by initializing the skill library with strategies derived from the expert-level dataset.
@@
Moreover, building autonomous agents to cooperate in completing tasks under partial observation requires an efficient communication protocol. However, naive communication leads to the risks of hallucination caused by meaningless chatter between agents (Li et al., 2024a). Inspired by entity-based MARL (Iqbal et al., 2021; Ding et al., 2023), we present a structured communication protocol to formulate the communication among agents along with a global memory that allows all agents to retrieve. The protocol incorporates a multi-hop propagation mechanism, enabling agents to infer information about entities beyond their field of view through information shared by teammates. Similar to previous approaches, each agent maintains a local memory to preserve current and historical experiences.
@@
Empirically, COMPASS demonstrates effective adaptation and skill synthesis in cooperative multi-agent scenarios. Through its dynamic skill library, it creates reusable and interpretable code-based behaviors that evolve during task execution. We evaluate COMPASS systematically in the improved StarCraft Multi-Agent Challenge (SMACv2) using both open-source (Qwen2-VL-72B (Wang et al., 2024c)) and closed-source (GPT-4o-mini [^1] https://platform.openai.com/docs/models#gpt-4o-mini, Claude-3-Haiku [^2] https://www.anthropic.com/news/claude-3-haiku) VLMs. COMPASS achieves strong results in Protoss scenarios with a 57% win rate, substantially outperforming state-of-the-art MARL algorithms, including QMIX (Rashid et al., 2020), MAPPO (Yu et al., 2022), HAPPO (Kuba et al., 2022), and HASAC (Liu et al., 2024). COMPASS maintains moderate performance in Terran scenarios and handles asymmetric settings effectively, though showing limited success in Zerg task. We further demonstrate COMPASS’s ability to bootstrap effective strategies from expert demonstrations.
@@ 2.1 Agents in the StarCraft Multi-Agent Challenge
SMAC (Samvelyan et al., 2019), a predominant cooperative MARL benchmark based on the StarCraft II real-time strategy game (Vinyals et al., 2017), focuses on decentralized micromanagement scenarios where each unit operates under decentralized execution with partial observability to defeat enemy units controlled by Starcraft II’s built-in AI opponent. Previous research in SMAC resort to MARL which can be divided into two categories: 1) Online MARL: One line of representative research is value decomposition (VD) (Rashid et al., 2020; Wang et al., 2021; Son et al., 2019), which decomposes the centralized action-value function into individual utility functions. On the other hand, multi-agent policy gradient (MAPG) methods (Yu et al., 2022; Kuba et al., 2022; Liu et al., 2024; Li et al., 2023; Wen et al., 2022; Hu et al., 2024; Na et al., 2024) extend single-agent policy gradient algorithm to multi-agent with coordination modeling. Researches such as MAPPO (Yu et al., 2022), HAPPO (Kuba et al., 2022), and HASAC (Liu et al., 2024) combine trust region and maximum entropy with MARL in a non-trivial way respectively. To encourage coordination, communication methods (Hu et al., 2024; Lo et al., 2024), sequential modeling methods (Li et al., 2023; Wen et al., 2022), and cooperative exploration methods (Mahajan et al., 2019; Na et al., 2024) have been proposed. 2) Offline MARL: Recent efforts such as MADT (Meng et al., 2023), ODIS (Zhang et al., 2022), and MADiff (Zhu et al., 2025) leverage data-driven training via pre-collected offline datasets to enhance policy training efficiency. However, the near-optimal performance of these existing approaches on SMAC highlights the benchmark’s limited stochasticity and partial observability. To address these limitations, SMACv2 (Ellis et al., 2024) introduces more complexity to necessitate decentralized closed-loop control policies. There have been some recent attempts (Li et al., 2024b; McClellan et al., 2024; Formanek et al., 2023; Li et al., 2024c) to evaluate MARL algorithms on SMACv2, and the results confirm the complexity. However, current learning-based multi-agent methods are computationally inefficient and non-interpretable. In the quest to find methods that are sample-efficient and interpretable, LLM-SMAC (Deng et al., 2024) leverage LLMs to generate centralized decision tree code under global information in an open-loop framework. Unlike prior works, COMPASS integrates a Vision-Language Model (VLM) with each agent in a decentralized closed-loop manner under partial observability, improving both real-world applicability and scalability.
@@ 2.2 LLM-based Multi-Agent System
Based on the inspiring capabilities of LLMs, such as zero-shot planning and complex reasoning (Kojima et al., 2024; Zhao et al., 2024; Wei et al., 2024; Besta et al., 2024), embodied single-agent researches have demonstrated the effectiveness of LLMs in solving complex long-horizon tasks (Wang et al., 2024b; Yao et al., 2023; Shinn et al., 2023; brian ichter et al., 2022; Ma et al., 2024b; Wang et al., 2024d; Tan et al., 2024). Despite significant advances in single-agent applications, developing real-world multi-agent systems with foundation models remains challenging, primarily due to the nature of decentralized control under partial observability in multi-agent settings (Pajarinen & Peltonen, 2011). Most prior efforts (Mandi et al., 2024; Zhang et al., 2024, 2025; Nayak et al., 2025; Gong et al., 2024) leverage a hierarchical framework with components like perception, communication, planning, execution, and memory to build multi-agent systems with collective behaviors. These approaches can be roughly classified into two groups. 1) Centralized plan: MindAgent (Gong et al., 2024) adopts a centralized planning scheme with a pre-defined oracle in a fully observable multi-agent game. LLaMAR (Nayak et al., 2025) employs LLMs to manage long-horizon tasks in partially observable environments without assumptions about access to perfect low-level policies. 2) Decentralized plan: ProAgent (Zhang et al., 2025) introduces Theory of Mind (ToM), enabling agents to reason about others’ mental states. RoCo (Mandi et al., 2024) and CoELA (Zhang et al., 2024) assign separate LLMs to each embodied agent for collaboration with communication. However, RoCo and CoELA assume a skill library with a low-level heuristic controller, which is impractical in real-world applications. Moreover, RoCo’s open-loop plan-and-execute paradigm fails to incorporate environmental feedback during decision-making. In contrast, our work does not assume any pre-defined low-level controller and generates code-based action through VLMs in a closed-loop manner.
@@ 3 Preliminaries
We model a fully cooperative multi-agent game with $N$ agents as a decentralized partially observable Markov decision process (Dec-POMDP) (Oliehoek & Amato, 2016), which is formally defined as a tuple $\mathcal{G}=(\mathcal{N},\mathcal{S},\mathcal{O},\mathbb{O},\mathcal{B},\mathcal{A},\mathcal{T},\Omega,R,\gamma,\rho_{0})$. $\mathcal{N}=\{1,\ldots,N\}$ is a set of agents, $s\in\mathcal{S}$ denotes the state of the environment and $\rho_{0}$ is the distribution of the initial state. $\mathcal{A}=\prod_{i=1}^{N}A^{i}$ is the joint action space, $\mathbb{O}=\prod_{i=1}^{N}O^{i}$ is the set of joint observations. At time step $t$, each agent $i$ receives an individual partial observation $o^{i}_{t}\in O^{i}$ given by the observation function $\mathcal{O}:(a_{t},s_{t+1})\mapsto P(o_{t+1}|a_{t},s_{t+1})$ where $a_{t},s_{t+1}$ and $o_{t+1}$ are the joint actions, states and joint observations respectively. Each agent $i$ uses a stochastic policy $\pi^{i}(a^{i}_{t}|h^{i}_{t},\omega^{i}_{t})$ conditioned on its action-observation history $h^{i}_{t}=(o^{i}_{0},a^{i}_{0},\ldots,o^{i}_{t-1},a^{i}_{t-1})$ and a random seed $\omega^{i}_{t}\in\Omega_{t}$ to choose an action $a^{i}_{t}\in A^{i}$. A belief state $b_{t}$ is a probability distribution over states at time $t$, where $b_{t}\in\mathcal{B}$, and $\mathcal{B}$ is the space of all probability distributions over the state space. Actions $a_{t}$ drawn from joint policy $\pi(a_{t}|s_{t},\omega_{t})$ conditioned on state $s_{t}$ and joint random seed $\omega_{t}=(\omega^{1}_{t},\ldots,\omega^{N}_{t})$ change the state according to transition function $\mathcal{T}:(s_{t},a^{1}_{t},\ldots,a^{N}_{t})\mapsto P(s_{t+1}|s_{t},a^{1}_{t},\ldots,a^{N}_{t})$. All agents share the same reward $r_{t}=R(s_{t},a^{1}_{t},\ldots,a^{N}_{t})$ based on $s_{t}$ and $a_{t}$. $\gamma$ is the discount factor for future rewards. Agents try to maximize the expected total reward, $\mathcal{J}(\pi)=\mathbb{E}_{s_{0},a_{0},\dots}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$, where $s_{0}\sim\rho_{0}(s_{0}),a_{t}\sim\pi(a_{t}|s_{t},\omega_{t})$.

### v2

@@ Abstract
Despite much progress in training distributed artificial intelligence (AI), building cooperative multi-agent systems with multi-agent reinforcement learning (MARL) faces challenges in sample efficiency, interpretability, and transferability. Unlike traditional learning-based methods that require extensive interaction with the environment, large language models (LLMs) demonstrate remarkable capabilities in zero-shot planning and complex reasoning. However, existing LLM-based approaches heavily rely on text-based observations and struggle with the non-Markovian nature of multi-agent interactions under partial observability. We present COMPASS, a novel multi-agent architecture that integrates vision-language models (VLMs) with a dynamic skill library and structured communication for decentralized closed-loop decision-making. The skill library, bootstrapped from demonstrations, evolves via planner-guided tasks to enable adaptive strategies. COMPASS propagates entity information through multi-hop communication under partial observability. Evaluations on the improved StarCraft Multi-Agent Challenge (SMACv2) demonstrate COMPASS’s strong performance against state-of-the-art MARL baselines across both symmetric and asymmetric scenarios. Notably, in the symmetric Protoss 5v5 task, COMPASS achieved a 57% win rate, representing a 30 percentage point advantage over QMIX (27%). Project page can be found at https://stellar-entremet-1720bb.netlify.app/.
@@ 1 Introduction
A major long-term goal for the field of cooperative multi-agent systems (MAS), e.g. multi-robot control Gao et al. (2024); Feng et al. (2024), power management Monroc et al. (2024), and multi-agent games Samvelyan et al. (2019); Kurach et al. (2020), is to build a protocol of collaboration among the agents. Multi-agent reinforcement learning (MARL), proven as an advanced paradigm of distributed artificial intelligence (AI), holds promise for discovering collective behavior from interactions. One line of works follows centralized training decentralized execution (CTDE) paradigm Rashid et al. (2020); Yu et al. (2022); Liu et al. (2024); Lowe et al. (2017); Zhu et al. (2025); Li et al. (2024b). CTDE assumes a central controller that exploits global information, while the individual policies are designed to allow for decentralized execution. However, in many real-world scenarios, the central controller becomes unfeasible due to the communication overhead that exponentially scales with the number of agents, thereby compromising the scalability of MAS. In contrast, the decentralized training decentralized execution paradigm (DTDE) discards this assumption and is scalable to large-scale systems Su & Lu (2024); Zhang et al. (2018); Ma et al. (2024a). However, DTDE requires complicated learning and planning under uncertainty, as partial observability magnifies the discrepancy between each agent’s local observation and global information. Although much progress has been made, MARL suffers from compromised sample efficiency, interpretability, and transferability.
@@
The emergence of Large Language Models (LLMs) has revitalized this field. LLM-based multi-agents have been proposed to leverage their remarkable capacity to perform task-oriented collective behaviors Mandi et al. (2024); Zhang et al. (2024); Gong et al. (2024); Zhang et al. (2025); Nayak et al. (2025). The LLMs are used for high-level planning to generate centralized Gong et al. (2024); Nayak et al. (2025); Deng et al. (2024) or decentralized plans Mandi et al. (2024); Zhang et al. (2024), often adopting a hierarchical decision-making structure in conjunction with a pre-defined low-level controller. While these methods have succeeded in a set of multi-agent problems including Overcooked-AI Carroll et al. (2019), SMAC Samvelyan et al. (2019), and VirtualHome Puig et al. (2018), heavy reliance on text-based observation prevents them from learning from multi-modal information. Moreover, they ignore the non-Markovian nature of MAS, where learning and planning necessitate a decentralized closed-loop solution.
@@
In this paper, we mitigate the existing research limitations and advance general decision-making for cooperative multi-agent systems. At a high level, COMPASS combines a vision language models (VLMs)-based planner with a dynamic skill library for storing and retrieving complex behaviors, along with a structured communication protocol. A diagram of COMPASS is provided in Figure. 1. Inspired by Cradle Tan et al. (2024), the VLM-based planner perceives the visual and textual observation and suggests the most suitable executable code from the skill library. We adopt the code-as-policy paradigm Wang et al. (2024d) instead of task-specific primitive actions, as it constrains generalizability and fails to fully leverage foundation models’ extensive world knowledge and sophisticated reasoning capabilities.
@@
Traditional open-loop methods struggle to produce effective plans that adapt to dynamics in stochastic, partially observable environments. To address this challenge, the VLM-based planer attempts to solve challenging and ambiguous final tasks, such as "Defeat all enemy units in the StarCraft multi-agent combat scenario while coordinating with allied units", by progressively proposing a sequence of clear, manageable sub-tasks while incorporating environmental feedback and task progress. COMPASS generates Python scripts through LLMs as semantic-level skills to accomplish sub-tasks, incrementally building a skill library throughout the task progress. Each skill is indexed through its documentation embeddings, enabling retrieval based on task-skill relevance. However, developing the skill library from scratch requires extensive exploration to discover viable strategies. In contrast, we pre-collect demonstration videos and introduce a "warm start" by initializing the skill library with strategies derived from the expert-level dataset.
@@
Moreover, building autonomous agents to cooperate in completing tasks under partial observation requires an efficient communication protocol. However, naive communication leads to the risks of hallucination caused by meaningless chatter between agents Li et al. (2024a). Inspired by entity-based MARL Iqbal et al. (2021); Ding et al. (2023), we present a structured communication protocol to formulate the communication among agents along with a global memory that allows all agents to retrieve. The protocol incorporates a multi-hop propagation mechanism, enabling agents to infer information about entities beyond their field of view through information shared by teammates. Similar to previous approaches, each agent maintains a local memory to preserve current and historical experiences.
@@
Empirically, COMPASS demonstrates effective adaptation and skill synthesis in cooperative multi-agent scenarios. Through its dynamic skill library, it creates reusable and interpretable code-based behaviors that evolve during task execution. We evaluate COMPASS systematically in the improved StarCraft Multi-Agent Challenge (SMACv2) using both open-source (Qwen2-VL-72B Wang et al. (2024c)) and closed-source (GPT-4o-mini [^1] https://platform.openai.com/docs/models#gpt-4o-mini, Claude-3-Haiku [^2] https://www.anthropic.com/news/claude-3-haiku) VLMs. COMPASS achieves strong results in Protoss scenarios with a 57% win rate, substantially outperforming state-of-the-art MARL algorithms, including QMIX Rashid et al. (2020), MAPPO Yu et al. (2022), HAPPO Kuba et al. (2022), and HASAC Liu et al. (2024). COMPASS maintains moderate performance in Terran scenarios and handles asymmetric settings effectively, though showing limited success in Zerg task. We evaluate the contribution of individual COMPASS components to its overall performance. We further demonstrate COMPASS’s ability to bootstrap effective strategies from expert demonstrations.
@@ 2.1 Agents in the StarCraft Multi-Agent Challenge
SMAC Samvelyan et al. (2019), a predominant cooperative MARL benchmark based on the StarCraft II real-time strategy game Vinyals et al. (2017), focuses on decentralized micromanagement scenarios where each unit operates under decentralized execution with partial observability to defeat enemy units controlled by Starcraft II’s built-in AI opponent. Previous research in SMAC resort to MARL which can be divided into two categories: 1) Online MARL: One line of representative research is value decomposition (VD) Rashid et al. (2020); Wang et al. (2021); Son et al. (2019), which decomposes the centralized action-value function into individual utility functions. On the other hand, multi-agent policy gradient (MAPG) methods Yu et al. (2022); Kuba et al. (2022); Liu et al. (2024); Li et al. (2023); Wen et al. (2022); Hu et al. (2024); Na et al. (2024) extend single-agent policy gradient algorithm to multi-agent with coordination modeling. Researches such as MAPPO Yu et al. (2022), HAPPO Kuba et al. (2022), and HASAC Liu et al. (2024) combine trust region and maximum entropy with MARL in a non-trivial way respectively. To encourage coordination, communication methods Hu et al. (2024); Lo et al. (2024), sequential modeling methods Li et al. (2023); Wen et al. (2022), and cooperative exploration methods Mahajan et al. (2019); Na et al. (2024) have been proposed. 2) Offline MARL: Recent efforts such as MADT Meng et al. (2023), ODIS Zhang et al. (2022), and MADiff Zhu et al. (2025) leverage data-driven training via pre-collected offline datasets to enhance policy training efficiency. However, the near-optimal performance of these existing approaches on SMAC highlights the benchmark’s limited stochasticity and partial observability. To address these limitations, SMACv2 Ellis et al. (2024) introduces more complexity to necessitate decentralized closed-loop control policies. There have been some recent attempts Li et al. (2024b); McClellan et al. (2024); Formanek et al. (2023); Li et al. (2024c) to evaluate MARL algorithms on SMACv2, and the results confirm the complexity. However, current learning-based multi-agent methods are computationally inefficient and non-interpretable. In the quest to find methods that are sample-efficient and interpretable, LLM-SMAC Deng et al. (2024) leverage LLMs to generate centralized decision tree code under global information in an open-loop framework. Unlike prior works, COMPASS integrates a Vision-Language Model (VLM) with each agent in a decentralized closed-loop manner under partial observability, improving both real-world applicability and scalability.
@@ 2.2 LLM-based Multi-Agent System
Based on the inspiring capabilities of LLMs, such as zero-shot planning and complex reasoning Kojima et al. (2024); Zhao et al. (2024); Wei et al. (2024); Besta et al. (2024), embodied single-agent researches have demonstrated the effectiveness of LLMs in solving complex long-horizon tasks Wang et al. (2024b); Yao et al. (2023); Shinn et al. (2023); brian ichter et al. (2022); Ma et al. (2024b); Wang et al. (2024d); Tan et al. (2024). Despite significant advances in single-agent applications, developing real-world multi-agent systems with foundation models remains challenging, primarily due to the nature of decentralized control under partial observability in multi-agent settings Pajarinen & Peltonen (2011). Most prior efforts Mandi et al. (2024); Zhang et al. (2024, 2025); Nayak et al. (2025); Gong et al. (2024) leverage a hierarchical framework with components like perception, communication, planning, execution, and memory to build multi-agent systems with collective behaviors. These approaches can be roughly classified into two groups. 1) Centralized plan: MindAgent Gong et al. (2024) adopts a centralized planning scheme with a pre-defined oracle in a fully observable multi-agent game. LLaMAR Nayak et al. (2025) employs LLMs to manage long-horizon tasks in partially observable environments without assumptions about access to perfect low-level policies. 2) Decentralized plan: ProAgent Zhang et al. (2025) introduces Theory of Mind (ToM), enabling agents to reason about others’ mental states. RoCo Mandi et al. (2024) and CoELA Zhang et al. (2024) assign separate LLMs to each embodied agent for collaboration with communication. However, RoCo and CoELA assume a skill library with a low-level heuristic controller, which is impractical in real-world applications. Moreover, RoCo’s open-loop plan-and-execute paradigm fails to incorporate environmental feedback during decision-making. In contrast, our work does not assume any pre-defined low-level controller and generates code-based action through VLMs in a closed-loop manner.
@@ 3 Preliminaries
We model a fully cooperative multi-agent game with $N$ agents as a decentralized partially observable Markov decision process (Dec-POMDP) Oliehoek & Amato (2016), which is formally defined as a tuple $\mathcal{G}=(\mathcal{N},\mathcal{S},\mathcal{O},\mathbb{O},\mathcal{B},\mathcal{A},\mathcal{T},\Omega,R,\gamma,\rho_{0})$. $\mathcal{N}=\{1,\ldots,N\}$ is a set of agents, $s\in\mathcal{S}$ denotes the state of the environment and $\rho_{0}$ is the distribution of the initial state. $\mathcal{A}=\prod_{i=1}^{N}A^{i}$ is the joint action space, $\mathbb{O}=\prod_{i=1}^{N}O^{i}$ is the set of joint observations. At time step $t$, each agent $i$ receives an individual partial observation $o^{i}_{t}\in O^{i}$ given by the observation function $\mathcal{O}:(a_{t},s_{t+1})\mapsto P(o_{t+1}|a_{t},s_{t+1})$ where $a_{t},s_{t+1}$ and $o_{t+1}$ are the joint actions, states and joint observations respectively. Each agent $i$ uses a stochastic policy $\pi^{i}(a^{i}_{t}|h^{i}_{t},\omega^{i}_{t})$ conditioned on its action-observation history $h^{i}_{t}=(o^{i}_{0},a^{i}_{0},\ldots,o^{i}_{t-1},a^{i}_{t-1})$ and a random seed $\omega^{i}_{t}\in\Omega_{t}$ to choose an action $a^{i}_{t}\in A^{i}$. A belief state $b_{t}$ is a probability distribution over states at time $t$, where $b_{t}\in\mathcal{B}$, and $\mathcal{B}$ is the space of all probability distributions over the state space. Actions $a_{t}$ drawn from joint policy $\pi(a_{t}|s_{t},\omega_{t})$ conditioned on state $s_{t}$ and joint random seed $\omega_{t}=(\omega^{1}_{t},\ldots,\omega^{N}_{t})$ change the state according to transition function $\mathcal{T}:(s_{t},a^{1}_{t},\ldots,a^{N}_{t})\mapsto P(s_{t+1}|s_{t},a^{1}_{t},\ldots,a^{N}_{t})$. All agents share the same reward $r_{t}=R(s_{t},a^{1}_{t},\ldots,a^{N}_{t})$ based on $s_{t}$ and $a_{t}$. $\gamma$ is the discount factor for future rewards. Agents try to maximize the expected total reward, $\mathcal{J}(\pi)=\mathbb{E}_{s_{0},a_{0},\dots}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$, where $s_{0}\sim\rho_{0}(s_{0}),a_{t}\sim\pi(a_{t}|s_{t},\omega_{t})$.